### v1

@@
[…]
@@
The question of whether video generative models can understand physical principles has been explored through a range of benchmarks designed to evaluate physical reasoning. Physion (17) and its successor Physion++ (18) use object collisions and stability to assess a model’s ability to predict physical outcomes and infer relevant properties of objects (e.g., mass, friction) during dynamic interactions. Similarly, CRAFT (19) and IntPhys (20) assess causal reasoning and intuitive physics, testing whether models can infer forces or understand object permanence. Intuitive physics has a rich history in Cognitive Science and is concerned with understanding how humans build a commonsense intuition for physical principles (e.g. 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31). Recent efforts have extended physical reasoning evaluation to generative video models. VideoPhy (32) and PhyGenBench (33) focus on assessing physical commonsense through text-based descriptions rather than visual data. These works emphasize logical reasoning about physical principles but do not incorporate real-world videos or dynamic visual contexts. The Cosmos project (34) aims to enable better embodied AI, including robotics. LLMPhy (35) combines a large language model with a non-differentiable physics simulator to iteratively estimate physical hyperparameters (e.g., friction, damping, layout) and predict scene dynamics. Other benchmarks, such as CoPhy (36) and CLEVERER (37), emphasize counterfactual reasoning and causal inference in video-based scenarios. ESPRIT (38) couples physical reasoning tasks with explainability via natural language explanations, and PhyWorld (39) evaluates the ability of generative video models to encode physical laws, focusing on physical realism.
Figure 3: A qualitative overview of recent synthetic datasets related to physical understanding (19, 20, 17, 38, 18, 36, 37, 39). These datasets are great for the purposes they were designed for, but not ideal for evaluating models trained on real-world videos due to the distribution shift.
However, a major drawback of these benchmarks is that the data they use is synthetic (see Fig 3 for samples). This introduces a real-vs-synthetic distribution shift that may confound results when testing video models trained on natural videos. The Physics-IQ dataset addresses this limitation by providing real-world videos, capturing diverse and complex physical phenomena (see Fig 1). With three views per scenario, controlled and measured physical variance (by recording two takes for each video), and challenging out-of-distribution settings it provides a rigorous design for evaluating video generative models.
@@
Video generative models commonly use metrics for evaluating the visual quality and realism of the generated videos. These metrics include Peak Signal-to-Noise Ratio (PSNR) (41), Structural Similarity Index Measure (SSIM) (42), Fréchet Video Distance (FVD) (43), and Learned Perceptual Image Patch Similarity (LPIPS) (44). These metrics are useful for comparing the appearance, temporal smoothness, and statistics of generated videos with the ground truth. Unfortunately, these metrics are not equipped to assess the understanding of physical laws by video models. For instance, both PSNR and SSIM evaluate pixel-level similarities but are not sensitive to the correctness of motion and interactions in a video; FVD captures overall feature distributions but does not penalize a model for physically implausible actions and LPIPS focuses on human-like perception of similarity rather than physical plausibility. While these metrics are great for measuring what they were designed for, they are not equipped to judge whether a model understands real-world physics.
@@
Figure 4: How well do current video generative models understand physical principles? Left. The Physics-IQ score is an aggregated measure across four individual metrics, normalized such that pairs of real videos that differ only by physical randomness score 100%. All evaluated models show a large gap, with the best model scoring 24.1%, indicating that physical understanding is severely limited. Right. In addition, the mean rank of models across all four metrics is shown here; the Spearman correlation between aggregated results on the left and mean rank on the right is high (-$\text{-}.87,\emph{p}<.005$), thus aggregating to a single Physics-IQ score largely preserves model rankings.
@@ Results
Figure 7: We here visualize success and failure scenarios within the fluid dynamics and solid mechanics categories for the two best models, VideoPoet and Runway Gen 3, according to our metrics. Both models are able to generate physics plausible frames for scenarios such as smearing paint on glass (VideoPoet) and pouring red liquid on a rubber duck (Runway Gen 3). At the same time, the models fail to simulate a ball falling into a crate or cutting a tangerine with a knife. See here for an animated version.
@@
The goal of our Physics-IQ benchmark is to understand, and quantify, whether generative video models learn physical principles. Therefore, we test all eight models in our study on every scenario and for each camera position (left, center, right) in the benchmark dataset. These samples are visualized in Figure 1. We first report the aggregated Physics-IQ results across all metrics related to physical understanding ($\mathsf{Spatial}\text{-}\mathsf{IoU},\penalty 10000\ \mathsf{Spatiotemporal}\text{-}\mathsf{IoU},\penalty 10000\ \mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU},\penalty 10000\ \mathsf{MSE}$) in Figure 4. The main takeaway from the left part of this figure is that all the models show a massive gap to the physical variance baseline, with the best model scoring only 24.1% out of the possible 100.0%. As we mentioned in the previous section, each scenario was recorded twice (take 1 and take 2) to estimate the natural variability in real-world physical phenomena. This estimate is termed the physical variance; the figure is normalized such that pairs of real videos that differ only by physical randomness score 100.0%. The gap between model performance and real videos demonstrates a severe lack of physical understanding in current powerful video generative models. Across the different models, VideoPoet (multiframe) (13) ranks best; interestingly, VideoPoet is a causal model. For the two models that have both an image2video (i2v) and a version conditioned on multiple frames (multiframe), the multiframe variants do better than the i2v variants. This is to be expected given that in order to predict the future as we require models to do on our challenging Physics-IQ benchmark, having access to temporal information (as multiframe variants have) should help.
@@
The performance of each model on each individual metric is reported in Table 1. As expected based on the aggregated results, VideoPoet (multiframe) performs best on a majority of the metrics (3 out of 4), while Stable Video Diffusion performs best on $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ followed by VideoPoet. The fact that Stable Video Diffusion does well in terms of producing action, while struggling with correct object placement and dynamics (as evidenced by weaker $\mathsf{Spatial}\text{-}\mathsf{IoU}$ and $\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$ scores) is also qualitatively evident in the generated videos which show more than usual movement, with the movement frequently being physically implausible and exaggerated. Similarly, Sora is among the best performing models for $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ but performs weaker on other metrics. Qualitatively, the generated videos from Sora are often visually and artistically superior, but they also frequently include transition cuts—despite instructed not to change the camera perspective—which is penalized by several other metrics. We expect that if a future version of this model more closely follows the prompt (static camera perspective, no camera movement), its Physics-IQ score would improve substantially. Qualitatively, the best and worst performing scenarios for the highest and lowest ranked models are visualized in Figure 7. These scenarios highlight both instances of physical understanding and failure cases.
@@
Our results indicate a striking lack of physical understanding in current generative video models. Why, then, do samples from many of those models that are circulated online look so realistic? We decided to quantify the visual realism of model-generated videos by asking a capable multimodal large language model, Gemini 1.5 Pro (45), to identify the generated one out of a pair of two videos for each scenario in the Physics-IQ benchmark). The result of this experiment is presented in Figure 5 (left). The MLLM score evaluates a model’s ability to generate videos that can deceive a multimodal large language model into classifying them as real. Accuracies that are closer to chance performance (50% by randomly guessing) indicate that the MLLM finds it harder to tell apart real from generated videos, thus attesting to the visual realism of the generated video. Overall, the MLLM frequently succeeds in identifying the model-generated video (up to 86.9% accuracy for Lumiere multiframe). However, one model stands out: Sora achieved the best MLLM score of 55.6%, outperforming all other models by a significant margin. Runway Gen 3 and VideoPoet (multiframe) ranked second and third, with scores of 74.8% and 77.3%, respectively, albeit with a considerable gap behind Sora. This finding quantifies that highly capable models such as Sora indeed succeed in generating visually realistic videos—even though their ability to understand physical principles, as shown in the previous section, is very limited. In fact, there is no significant correlation between visual realism and physical understanding: Figure 5 (right) demonstrates that there is a distinction between generating realistic videos and comprehending the physical principles of reality.
@@ Discussion
It may be worth pointing out that even though the models in our study often failed to generate physically plausible continuations, most current models were already successful on some scenarios. For example, the highest-ranking model, VideoPoet (multiframe), displayed remarkable physical understanding in certain scenarios—such as accurately simulating paint smearing on glass. In contrast, many lower-ranking models exhibited fundamental errors, such as physically implausible interactions (e.g., objects passing through other objects). Separate studies (e.g. 39) based on synthetic datasets have shown that given a large enough dataset size, video models are able to learn specific physical laws. We consider it likely that as models are trained on larger and more diverse corpora of videos, their understanding of real-world physics will continue to improve. We hope that quantifying physical understanding, as we aim to do by open-sourcing the physics-IQ benchmark, might facilitate tracking progress in this area.
@@
We observed a striking discrepancy between visual realism and physical understanding: those two properties are not statistically significantly correlated (Figure 5), thus models that produce the most realistic-looking videos don’t necessarily show the most physically-plausible continuations. For instance, in a scenario where a burning matchstick is lowered into a glass full of water (leading to the flame being extinguished), Runway Gen 3 generates a continuation where as soon as the flame touches the water, a candle spontaneously appears and is lit by the match. Every single frame of the video is high quality in terms of resolution and realism, but the temporal sequence is physically impossible. Such a tendency to hallucinate objects into existence is a drawback of many generative models and an active area of research in deep learning (47). In our experiments, we observed hallucinations in all models, but more powerful models like Runway Gen 3 and Sora often hallucinated information that was at least consistent with the scenario (e.g. matchstick - candle) rather than arbitrary, indicating at least a certain level of understanding.
@@
We intentionally designed Physics-IQ to be a challenging benchmark in order to provide useful signal for model development in the future; in this context it may be worth noting that our metrics may be on the conservative side by strongly penalizing object hallucinations, camera movement (which we prompted models not to do) or shot changes. For instance, Sora tends to show these more frequently than other models, leading to low scores on some metrics. This is not ideal, but we believe that in an area like deep learning where hype is omnipresent, scientific benchmarks should rather err on the side of caution.
Outlook.
The obvious trends are better models trained on bigger datasets. Will these models essentially solve physical understanding—or instead, hit a limit after which one can only improve one’s understanding the world by interacting with it? The jury is still out on this question, but the benchmark and analyses introduced in this article might help quantifying this either way. In addition to future models, improvements could also come from inference-time compute, such as scaling the number of samples. If this would lead to strong results, it would raise the following question: from a model’s perspective, is reality but one possibility among infinitely many others?
@@
[…]
@@
10:  $i\leftarrow\lfloor\alpha\rfloor$ $\triangleright$ Index of the first frame for interpolation
11:  $\beta\leftarrow\alpha-i$ $\triangleright$ Weight for linear interpolation
@@
[…]
@@
5:Initialize background model $\text{B}\leftarrow\frac{1}{w}\sum_{i=1}^{w}f_{i}$ $\triangleright$ Initial average reduces noise
@@
10:  Threshold $m_{t}\leftarrow 255$ if $d_{t}>\tau$, else 0

### v2

@@
Saman Motameda,1 Laura Culpb Kevin Swerskyb Priyank Jainib,$\dagger$ and Robert Geirhosb,$\dagger$
@@
The question of whether video generative models can understand physical principles has been explored through a range of benchmarks designed to evaluate physical reasoning. Physion (17) and its successor Physion++ (18) use object collisions and stability to assess a model’s ability to predict physical outcomes and infer relevant properties of objects (e.g., mass, friction) during dynamic interactions. Similarly, CRAFT (19) and IntPhys (20) assess causal reasoning and intuitive physics, testing whether models can infer forces or understand object permanence. Intuitive physics has a rich history in Cognitive Science and is concerned with understanding how humans build a commonsense intuition for physical principles (e.g. 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31). Recent efforts have extended physical reasoning evaluation to generative video models. VideoPhy (32) and PhyGenBench (33) focus on assessing physical commonsense through text-based descriptions rather than visual data. These works emphasize logical reasoning about physical principles but do not incorporate real-world videos or dynamic visual contexts. PhysGame (34) focuses on gameplay, while the Cosmos project (35) aims to enable better embodied AI, including robotics. LLMPhy (36) combines a large language model with a non-differentiable physics simulator to iteratively estimate physical hyperparameters (e.g., friction, damping, layout) and predict scene dynamics. Other benchmarks, such as CoPhy (37) and CLEVERER (38), emphasize counterfactual reasoning and causal inference in video-based scenarios. ESPRIT (39) couples physical reasoning tasks with explainability via natural language explanations, and PhyWorld (40) evaluates the ability of generative video models to encode physical laws, focusing on physical realism. A comprehensive overview of recent models and methods is provided by (41).
Figure 3: A qualitative overview of recent synthetic datasets related to physical understanding (19, 20, 17, 39, 18, 37, 38, 40). These datasets are great for the purposes they were designed for, but not ideal for evaluating models trained on real-world videos due to the distribution shift.
However, a major drawback of many benchmarks is that the data they use is synthetic (see Fig 3 for samples). This introduces a real-vs-synthetic distribution shift that may confound results when testing video models trained on natural videos. The Physics-IQ dataset overcomes this limitation by providing real-world videos, capturing diverse and complex physical phenomena (see Fig 1). With three views per scenario, controlled and measured physical variance (by recording two takes for each video), and challenging out-of-distribution settings it provides a rigorous design for evaluating video generative models.
@@
Video generative models commonly use metrics (43, 44, 45, 46) and benchmarks (47, 48, 49) suited for evaluating the visual quality and realism of the generated videos. These metrics include Peak Signal-to-Noise Ratio (PSNR) (43), Structural Similarity Index Measure (SSIM) (44), Fréchet Video Distance (FVD) (45, 50), and Learned Perceptual Image Patch Similarity (LPIPS) (46). These metrics are useful for comparing the appearance, temporal smoothness, and statistics of generated videos with the ground truth. Unfortunately, these metrics are not equipped to assess the understanding of physical laws by video models. For instance, both PSNR and SSIM evaluate pixel-level similarities but are not sensitive to the correctness of motion and interactions in a video; FVD captures overall feature distributions but does not penalize a model for physically implausible actions and LPIPS focuses on human-like perception of similarity rather than physical plausibility. While these metrics are great for measuring what they were designed for, they are not equipped to judge whether a model understands real-world physics.
@@
Figure 4: How well do current video generative models understand physical principles? Left. The Physics-IQ score is an aggregated measure across four individual metrics, normalized such that pairs of real videos that differ only by physical randomness score 100%. All evaluated models show a large gap, with the best model scoring 24.1%, indicating that physical understanding is severely limited. Right. In addition, the mean rank of models across all four metrics is shown here; the Spearman correlation between aggregated results on the left and mean rank on the right is high (-.$\text{-}.87,\emph{p}<.005- .$), thus aggregating to a single Physics-IQ score largely preserves model rankings.
@@ Results
Figure 7: We here visualize success and failure scenarios within the fluid dynamics and solid mechanics categories for the two best models, VideoPoet and Runway Gen 3, according to our metrics. Both models are able to generate physics plausible frames for scenarios such as smearing paint on glass (VideoPoet) and pouring red liquid on a rubber duck (Runway Gen 3). At the same time, the models fail to simulate a ball falling into a crate or cutting a tangerine with a knife. See here for an animated version of this figure.
@@
The goal of our Physics-IQ benchmark is to understand, and quantify, whether generative video models learn physical principles. Therefore, we test all eight models in our study on every scenario and for each camera position (left, center, right) in the benchmark dataset. These samples are visualized in Figure 1. We first report the aggregated Physics-IQ results across all metrics related to physical understanding ($\mathsf{Spatial}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{Spatiotemporal}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU},\leavevmode\nobreak\ \mathsf{MSE}$) in Figure 4. The main takeaway from the left part of this figure is that all the models show a massive gap to the physical variance baseline, with the best model scoring only 24.1% out of the possible 100.0%. As we mentioned in the previous section, each scenario was recorded twice (take 1 and take 2) to estimate the natural variability in real-world physical phenomena. This estimate is termed the physical variance; the figure is normalized such that pairs of real videos that differ only by physical randomness score 100.0%. The gap between model performance and real videos demonstrates a severe lack of physical understanding in current powerful video generative models. Across the different models, VideoPoet (multiframe) (13) ranks best; interestingly, VideoPoet is a causal model. For the two models that have both an image2video (i2v) and a version conditioned on multiple frames (multiframe), the multiframe variants do better than the i2v variants. This is to be expected given that in order to predict the future as we require models to do on our challenging Physics-IQ benchmark, having access to temporal information (as multiframe variants have) should help.
@@
The performance of each model on each individual metric is reported in Table 1. As expected based on the aggregated results, VideoPoet (multiframe) performs best on a majority of the metrics (3 out of 4), while Stable Video Diffusion performs best on $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ followed by VideoPoet. The fact that Stable Video Diffusion does well in terms of producing action, while struggling with correct object placement and dynamics (as evidenced by weaker $\mathsf{Spatial}\text{-}\mathsf{IoU}$ and $\mathsf{Spatiotemporal}\text{-}\mathsf{IoU}$ scores) is also qualitatively evident in the generated videos which show more than usual movement, with the movement frequently being physically implausible and exaggerated. Similarly, Sora is among the best performing models for $\mathsf{Weighted}\text{-}\mathsf{spatial}\text{-}\mathsf{IoU}$ but performs weaker on other metrics. Qualitatively, the generated videos from Sora are often visually and artistically superior, but they also frequently include transition cuts—despite instructed not to change the camera perspective—which is penalized by several other metrics. We expect that if a future version of this model more closely follows the prompt (static camera perspective, no camera movement), its Physics-IQ score would improve substantially. Qualitatively, success and failure cases are visualized in Figure 7.
@@
Our results indicate a striking lack of physical understanding in current generative video models. Why, then, do samples from many of those models that are circulated online look so realistic? We decided to quantify the visual realism of model-generated videos by asking a capable multimodal large language model, Gemini 1.5 Pro (51), to identify the generated one out of a pair of two videos for each scenario in the Physics-IQ benchmark). The result of this experiment is presented in Figure 5 (left). The MLLM score evaluates a model’s ability to generate videos that can deceive a multimodal large language model into classifying them as real. Accuracies that are closer to chance performance (50% by randomly guessing) indicate that the MLLM finds it harder to tell apart real from generated videos, thus attesting to the visual realism of the generated video. Overall, the MLLM frequently succeeds in identifying the model-generated video (up to 86.9% accuracy for Lumiere multiframe); that said, the model-produced explanations for decisions are often unrelated to the visual content, akin to post-hoc rationalizations known from human experiments (52). One model particularly stands out: Sora achieved the best MLLM score of 55.6%, outperforming all other models by a significant margin. Runway Gen 3 and VideoPoet (multiframe) ranked second and third, with scores of 74.8% and 77.3%, respectively, albeit with a considerable gap behind Sora. Thus, highly capable models such as Sora indeed succeed in generating visually realistic videos—even though their ability to understand physical principles, as shown in the previous section, is very limited. This finding aligns with a number of intriguing studies discovering that many deep learning models also lack an understanding of intuitive physics and causal reasoning (53, 54, 55, 56, 57, 58) and are sometimes described as “blind” (59). For video models, we find no significant correlation between visual realism and physical understanding: Figure 5 (right) demonstrates that there is a distinction between generating realistic videos and comprehending the physical principles of reality.
@@ Discussion
It may be worth pointing out that even though the models in our study often failed to generate physically plausible continuations, most current models were already successful on some scenarios. For example, the highest-ranking model, VideoPoet (multiframe), displayed remarkable physical understanding in certain scenarios—such as accurately simulating paint smearing on glass. In contrast, many lower-ranking models exhibited fundamental errors, such as physically implausible interactions (e.g., objects passing through other objects). Separate studies (e.g. 40) based on synthetic datasets have shown that given a large enough dataset size, video models are able to learn specific physical laws. We consider it likely that as models are trained on larger and more diverse corpora of videos, their understanding of real-world physics will continue to improve. We hope that quantifying physical understanding, as we aim to do by open-sourcing the physics-IQ benchmark, might facilitate tracking progress in this area.
@@
We observed a striking discrepancy between visual realism and physical understanding: those two properties are not statistically significantly correlated (Figure 5), thus models that produce the most realistic-looking videos don’t necessarily show the most physically-plausible continuations. For instance, in a scenario where a burning matchstick is lowered into a glass full of water (leading to the flame being extinguished), Runway Gen 3 generates a continuation where as soon as the flame touches the water, a candle spontaneously appears and is lit by the match. Every single frame of the video is high quality in terms of resolution and realism, but the temporal sequence is physically impossible. Such a tendency to hallucinate objects into existence is a drawback of many generative models and an active area of research in deep learning (61). In our experiments, we observed hallucinations in all models, but more powerful models like Runway Gen 3 and Sora often hallucinated information that was at least consistent with the scenario (e.g. matchstick - candle), indicating at least a certain level of understanding.
@@
We intentionally designed Physics-IQ to be a challenging benchmark in order to provide useful signal for model development in the future; in this context it may be worth noting that our metrics may be on the conservative side by strongly penalizing object hallucinations, camera movement (which we prompted models not to do) or shot changes. For instance, Sora tends to show these more frequently than other models, leading to low scores on some metrics. This is not ideal, but we believe that in an area like deep learning where hype is omnipresent, scientific benchmarks should err on the side of caution.
Outlook: Understanding without interaction?
Our findings are connected to a larger, interdisciplinary debate at the heart of intelligence: Does an understanding of the world emerge from predicting what happens next (next-video-frame prediction in artificial intelligence, predictive coding in neuroscience)—or, alternatively, is it necessary to interact with the world in order to understand it (as argued by proponents of embodied cognition and robotics)? In cognitive science, being able to interact with the world is seen as an important component for developing intuitive physics (62, 63, 64, 65), in combination with predicting the outcome of a person’s actions (66, 67, 68). In contrast, deep learning’s current bread-and-butter approach is scaling models and datasets without interactions. Will these models essentially solve physical understanding—or instead, hit a limit after which one can only improve one’s understanding of the world by interacting with it? The jury is still out on this question, but the benchmark and analyses introduced in this article might help quantifying this either way. In addition to future models, improvements could also come from inference-time scaling (69, 70, 71) such as sampling more. If this would indeed lead to strong results, it would raise the following question: from a model’s perspective, is reality but one possibility among infinitely many others?
@@
3:$f_{\mathsf{original}}\leftarrow$ extract frames from $V$ at $\mathsf{fps_{original}}$
4:$\mathsf{duration}\leftarrow$ length of of $V$
5:$n_{\mathsf{original}}\leftarrow$ number of frames in $\mathsf{fps_{original}}$
6:$n_{\mathsf{new}}\leftarrow\mathsf{duration}\cdot\mathsf{fps_{new}}$
@@
10:     $i\leftarrow\lfloor\alpha\rfloor$triangleright Index of the first frame for interpolation
11:     $\beta\leftarrow\alpha-i$triangleright Weight for linear interpolation
@@
18:$V^{\prime}\leftarrow$ recreate video from $f_{\mathsf{new}}$ with $\mathsf{fps_{new}}$
@@
5:Initialize background model $\text{B}\leftarrow\frac{1}{w}\sum_{i=1}^{w}f_{i}$triangleright Initial average reduces noise
@@
10:     Threshold $m_{t}\leftarrow 255$ if $d_{t}>\tau$, else 00